OpenStudy (anonymous):

A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (solomonzelman):

I would suggest to multiply the 1st equation times 3, and then subtract the equations . Alternatively, you can multiply the first equation times -3, and add the equations. Makes no difference.

OpenStudy (anonymous):

I really don't know how to do this, I'm just trying to finish up for the segment. :(

OpenStudy (solomonzelman):

lets do a multiplication times -3, and then adding, okay>

OpenStudy (solomonzelman):

So, when you multiply \(\large\color{black}{ -3x + 7y = -16 }\) times -3, you are multiplying EVERY THING in the equation times -3. So can you tell me: \(\large\color{black}{ ( -3x)\times (-3)=? }\) \(\large\color{black}{ ( 7y)~~~\times (-3)=? }\) \(\large\color{black}{ ( -16)\times (-3)=? }\)

OpenStudy (anonymous):

9 -21 and 48?

OpenStudy (solomonzelman):

you are forgetting the Ys and Xs for the first 2. For the last one, 48 is correct.

OpenStudy (anonymous):

But what are the values for x & y?

OpenStudy (solomonzelman):

when you multiply \(\large\color{black}{ ( -3y)\times (-3) }\) you get \(\large\color{black}{ 9y }\), not just \(\large\color{black}{ 9 }\) correct?

OpenStudy (solomonzelman):

I mean 3x, not 3y.

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

So what would I put for Part A? Sorry to rush its timed though :/

OpenStudy (solomonzelman):

for part A, you need to just add the equations, and this will be an equation 1 of your new system. and the second equation of your new system, is a multiply of any of the initial two equations.

OpenStudy (anonymous):

But how would I type it up? I'm really confused I'm sorry.

OpenStudy (solomonzelman):

okay, can you add your 2 equations? -3x + 7y = -16 -9x + 5y = 16 ------------

OpenStudy (anonymous):

Is the first one 36?

OpenStudy (anonymous):

I added the stuff from before..

OpenStudy (solomonzelman):

add the equations, please. I don't know what just a plain 36 means.

OpenStudy (anonymous):

x + y = 36??? I really don't know.

OpenStudy (solomonzelman):

-3x+(-9x)=? (7y)+(5y)=? (-16)+(16)=?

OpenStudy (anonymous):

-12x 12y and 0

OpenStudy (solomonzelman):

so you have: -12x + 12y = 0 as your first equation.

OpenStudy (solomonzelman):

then choose any equation out of -3x + 7y = -16 -9x + 5y = 16 and multiply it by any number (besides zero). I would just multiply the 1st equation times -2.

OpenStudy (anonymous):

Wait what?

OpenStudy (solomonzelman):

multiply any of your (initial) equations by any number. This is what they want, they want the second new equation to be a multiple of one of the first two.

OpenStudy (solomonzelman):

Like multiply the 1st equation times -2. -3x + 7y = -16 times(-2), (-3x)(-2) + (7y)(-2) = (-16)(-2) do you see what I am doing? what do you get?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

im sorry to say this but can you please just give me the answer for this one? I really can't figure this out and I'm running out of time and I still have to work on other questions.

OpenStudy (solomonzelman):

no you are multiplying everything as I said; (-3x)(-2) + (7y)(-2) = (-16)(-2) you just need to simplify. And sorry i lost connection.

OpenStudy (anonymous):

-5x + -14?

OpenStudy (anonymous):

-14y *

OpenStudy (solomonzelman):

not exactly: (-3x)(-2) + (7y)(-2) = (-16)(-2) 6x + (-14y) = 32 6x -14y = 32 \(\normalsize\color{blue}{ \rm correct ? }\)

OpenStudy (solomonzelman):

so there is your system that you need for A. -12x + 12y = 0 6x -14y = 32

OpenStudy (anonymous):

Oh oops! I added -3x and -2

OpenStudy (solomonzelman):

yes lol...

OpenStudy (anonymous):

So that's what I put for part a?

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

And what do I put for part b?

OpenStudy (solomonzelman):

This is the system you had initially. -3x + 7y = -16 -9x + 5y = 16 This is the system that you created. -12x + 12y = 0 6x -14y = 32 you are to show that these two systems have the same solution, by solving each one of them.

OpenStudy (anonymous):

So how would I lay that out?

OpenStudy (solomonzelman):

you just need to solve each system. let's start from the -3x + 7y = -16 -9x + 5y = 16 okay?

OpenStudy (anonymous):

Wasn't it -12x - 12y = 0?

OpenStudy (solomonzelman):

this is just the addition of the 2 equations of your intial system that you needed for the 1st equation of part A.

OpenStudy (solomonzelman):

You need to solve for x and y. -3x + 7y = -16 -9x + 5y = 16 multiply the 1st equation times -3. (we have already done that, I think)

OpenStudy (anonymous):

Yeah it was -48 right?

OpenStudy (solomonzelman):

you are multiplying every term by -3. the -3x times -3, 7y times -3, and -16 times -3.

OpenStudy (solomonzelman):

\(\large\color{black}{ (-3x)(-3) + (7y)(-3) = (-16)(-3) }\)

OpenStudy (anonymous):

9x -2y & 48

OpenStudy (anonymous):

-21y*

OpenStudy (solomonzelman):

when we did it before (in this same question) we got. \(\large\color{black}{ 9x-21y=48 }\) you are correct.

OpenStudy (solomonzelman):

you had: -3x + 7y = -16 -9x + 5y = 16 multiplied the first equation times -3 and now have: 9x -21y = 48 -9x + 5y = 16 now, please add the equations for me. 9x -21y = 48 -9x + 5y = 16 --------------

OpenStudy (anonymous):

64?

OpenStudy (solomonzelman):

9x + (-9x) =? -21y + 5y= ? and 64 (for the sum of 48 and 16) is correct.

OpenStudy (anonymous):

0 & -16y

OpenStudy (solomonzelman):

yes. so when you add you get. -16y = 64 (and as you can notice, the Xs cancel) can you solve for y?

OpenStudy (anonymous):

-4?

OpenStudy (solomonzelman):

yes, y=-4.

OpenStudy (solomonzelman):

now, plug in -4 instead of y into any of the equations of the original system -3x + 7y = -16 -9x + 5y = 16

OpenStudy (solomonzelman):

if you plug -4 for y into both of the equations and get the same answer (though), this will not only find the x, but also verify and check.

OpenStudy (anonymous):

Yeah I got the same answer

OpenStudy (anonymous):

so that's what I put?

OpenStudy (solomonzelman):

you got the same answer for both, what did you get for x?